# Anomaly-induced vanishing of brane partition functions

### v1

@@ Appendix B Twisted fiber integration
where the representation $\rho:\pi_{1}(B)\to\text{Aut}_{\mathbb{Z}}(H^{\ast}(F;G))$ is regarded as a $\mathbb{Z}[\pi_{1}(B)]$-module, and so $H^{\ast}(B;H^{\ast}(F;G)_{\rho})$ is taken to be cohomology with local coefficients MR1867354 . Importantly, the spectral sequence abuts to ordinary cohomology of the total space $X$.
@@ Appendix C Stiefel-Whitney classes of lens spaces
When $k=2m$ is even, $H^{i}(S^{2n+1}/\mathbb{Z}_{k};\mathbb{Z}_{2})=\mathbb{Z}_{2}$ and $H^{i}(S^{2n+1}/\mathbb{Z}_{k};\mathbb{Z}_{k})=\mathbb{Z}_{k}$ for all $0\leq i\leq 2n+1$. If we denote the generators $\hat{a}_{1}\in H^{1}(S^{2n+1}/\mathbb{Z}_{k};\mathbb{Z}_{k})$ and $\hat{b}_{2}\in H^{2}(S^{2n+1}/\mathbb{Z}_{k};\mathbb{Z}_{k})$, then it follows from the simplicial complex of the lens space that (cf. Examples 3.9 and 3.41 in MR1867354 )

### v3

@@ Appendix B Twisted fiber integration
where the representation $\rho:\pi_{1}(B)\to\text{Aut}_{\mathbb{Z}}(H^{\ast}(F;G))$ is regarded as a $\mathbb{Z}[\pi_{1}(B)]$-module, and so $H^{\ast}(B;H^{\ast}(F;G)_{\rho})$ is taken to be cohomology with local coefficients Hatcher (2002). Importantly, the spectral sequence abuts to ordinary cohomology of the total space $X$.
@@ Appendix C Stiefel-Whitney classes of lens spaces
When $k=2m$ is even, $H^{i}(S^{2n+1}/\mathbb{Z}_{k};\mathbb{Z}_{2})=\mathbb{Z}_{2}$ and $H^{i}(S^{2n+1}/\mathbb{Z}_{k};\mathbb{Z}_{k})=\mathbb{Z}_{k}$ for all $0\leq i\leq 2n+1$. If we denote the generators $\hat{a}_{1}\in H^{1}(S^{2n+1}/\mathbb{Z}_{k};\mathbb{Z}_{k})$ and $\hat{b}_{2}\in H^{2}(S^{2n+1}/\mathbb{Z}_{k};\mathbb{Z}_{k})$, then it follows from the simplicial complex of the lens space that (cf. Examples 3.9 and 3.41 in Hatcher (2002))